Week of 190429

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local: Miroslav (Chair, DB), Borja (Monitoring), Maarten (ALICE), Roberto (Storage)
  • remote: Christoph (CMS), Darren (RAL), Di (TRIUMF), Dave (FNAL), Jose (PIC), Ville (NDGF), Vincenzo (EGI)

Experiments round table:

  • CMS reports ( raw view) -
    • Good CPU utilization ~225k cores
      • As usual ~75% production, 25% Analysis
    • Some instabilities of CMS data bookkeeping system (DBS)
      • Under investigation
    • High rate of authentications led to EOS issues
      • Discussing with EOS experts how to mitigate (RQF:1288700)

  • ALICE -
    • No major issues
    • Prague back in business after 2 months of network issues

  • LHCb reports ( raw view) -
    • Activity
      • User jobs, MC productions, and staging this week
    • Issues
      • RAL:
        • Continuing migration from Castor to ECHO
      • IN2P3:
        • Unscheduled warning downtime this morning for Patch for NFS mount problem

Sites / Services round table:

  • ASGC: NC
  • BNL: NC
  • CNAF:
  • EGI: NTR
  • IN2P3:
    • Since last maintenance early April, we are facing instabilities with batch system due to problem with NFS mounting point for local shared storage space. Problems are under control but not fully resolved (waiting answers from provider support).
  • KIT: NC
  • NL-T1: NC
  • NRC-KI: NC
  • OSG: NC
  • PIC: NTR
  • RAL: NTR

  • CERN computing services: NTR
  • CERN storage services:
    • FTS: The 2 FTS WLCG production instances at CERN will be upgraded to FTS v3.8.5, today 29/04 from 14:00 to 17:00 CEST. (OTG:0050038). The intervention will be transparent to clients.
  • CERN databases: NTR
  • Monitoring:
    • Final reports for the March 2019 availability sent around
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR


Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2019-04-30 - MiroslavPotocky
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback